Based on historical data, your manager believes that 45% of the company's orders come from first-time customers. A random sample of 70 orders will be used to estimate the proportion of first-time-customers. What is the probability that the sample proportion is between 0.24 and 0.48?



Answer = (Enter your answer as a number accurate to 4 decimal places.)
Answer #1

Solution

Given that,

p = 0.45

1 - p = 1 - 0.45 = 0.55

n = 70

\mu\hat p = p = 0.45

\sigma\hat p  \sqrt[p ( 1 - p ) / n] = \sqrt [(0.45 * 0.55) / 70 ] = 0.0595

P( 0.24 < \hat p < 0.48 )

= P[(0.24 - 0.45) / 0.0595 < ( \hat p - \mu \hat p ) / \sigma \hat p < (0.48 - 0.45) / 0.0595]

= P(-3.53 < z < 0.50)

= P(z < 0.50) - P(z < -3.53)

Using z table,   

= 0.6915 - 0.0002

= 0.6913